OpenStudy (anonymous):

What is the value of the x variable in the solution to the following system of equations? 4x + 2y = 12 x – y = 3

OpenStudy (anonymous):

Write y from the second equation in terms of x, y=x-3 and plug it in the first equation. 4x+2(x-3)=12 => 4x+2x-6=12 => x=1. Now go back to y=x-3, plug x=1, and you get y=2. So your solution set is {(1,2)}.

OpenStudy (anonymous):

Choices....... 0 –3 3 2

OpenStudy (anonymous):

4x + 2y = 12 x – y = 3 so x=y+3 then substitute 4(y+3)+2y=12 y=0 then x=y+3 x=3

OpenStudy (anonymous):

0